Calculate the resistance and inductance given a frequency of 50hz, applied voltage of 240v, capacitance of 40*10^-6f and current of 1.5<-35
1
Expert's answer
2021-09-21T12:01:57-0400

"Resistance, R=V/A=240/1.5=160\Omega"

calculate inductance L

at resonance "X_L=X_C=R"

So "X_L=2\pi fL"

"R=2\pi fL"

"L=\frac{R}{2\pi f}=\frac{160}{2\times \pi \times 50}=0.5093H"